OpenStudy (anonymous):

The ratio of the width and the length of a rectangular lot is 3:5. If their sum is 96m, what is the area of the rectangular lot?

OpenStudy (anonymous):

\[w/l=3/5\] w+l=96

OpenStudy (anonymous):

Thank youuuuu.

OpenStudy (anonymous):

3x+5x=96 x=16 area=48*80 area=3840